OpenStudy (anonymous):

HELP.What is the radius of a circle with the equation...x^2+y^2+4x+8y-10=0...round your answer to the nearest thousandth

OpenStudy (anonymous):

sqrt(30) = 5.477

OpenStudy (anonymous):

Which of the following is a degenerate circle a-x^2+y^2=121 b-(x-5)^2+(y-1)^2=0 c-x+y=12 dx^2+y^2=-4

OpenStudy (anonymous):

B : degenerate circle : radius = 0

OpenStudy (anonymous):

The graph of a degenerate circle is a _____. ellipse,point,line,circle

OpenStudy (anonymous):

point

OpenStudy (anonymous):

Which of the following graphs matches the circle defined by this equation? (x+2)^2+(y-4)^2=9

OpenStudy (anonymous):

the one with center at (-2, 4) and a radius of 3

OpenStudy (anonymous):

Referral Denied You don't have permission to access " http://media.apexlearning.com/Images/200706/04/ffa6a68c-643c-4446-9da1-3a3dd08c7501.jpg " on this server.

OpenStudy (anonymous):

flutter,what do you suggest

OpenStudy (anonymous):

if you're familiar with image uploading you could use one of those sites

OpenStudy (anonymous):

like tinypic, imageshack,others

OpenStudy (anonymous):

i'm really not is there anyway you could tell me which points the circle goes through ?

OpenStudy (anonymous):

has center at (-2, 4) and a radius of 3

OpenStudy (anonymous):

alright i got it !

OpenStudy (anonymous):

Which equation describes the circle having center point (3,5) and radius r = 2 in standard form? a(x+3)^2+(y+5)^2=2 b(x-3)^2+(y-5)^2=4 c(x-3)^2+(y-5)^2 d(x+3)^2+(y+5)^2=4

OpenStudy (anonymous):

B

OpenStudy (anonymous):

this is the last one :)

OpenStudy (anonymous):

Thanks so much you were a very good help !

OpenStudy (anonymous):

glad i could help :) best of luck